solve by elimination method
5x+5y=-11
7x-3y=19

Answer by user_dude2008(1862)

5x+5y=-11
7x-3y=19


3(5x+5y)=3(-11)
5(7x-3y)=5(19)


15x+15y=-33
35x-15y=95


50x=62
x=62/50=31/25

5x+5y=-11
5(31/25)+5y=-11
31/5+5y=-11
31+25y=-55
25y=-55-31
25y=-86
y=-86/25

Answers: x=31/25, y=-86/25